# A Research Roadmap for Augmenting Software Engineering Processes and Software Products with Generative AI

### v1

@@ Abstract.
Abstract. Generative AI (GenAI) is rapidly transforming software engineering (SE) practices, influencing how SE processes are executed, as well as how software systems are developed, operated, and evolved. This paper applies design science research to build a roadmap for GenAI-augmented SE. The process consists of three cycles that incrementally integrate multiple sources of evidence, including collaborative discussions from the FSE 2025 “Software Engineering 2030” workshop, rapid literature reviews, and external feedback sessions involving peers. McLuhan’s tetrads were used as a conceptual instrument to systematically capture the transforming effects of GenAI on SE processes and software products. The resulting roadmap identifies four fundamental forms of GenAI augmentation in SE and systematically characterizes their related research challenges and opportunities. These insights are then consolidated into a set of future research directions. By grounding the roadmap in a rigorous multi-cycle process and cross-validating it among independent author teams and peers, the study provides a transparent and reproducible foundation for analyzing how GenAI affects SE processes, methods and tools, and for framing future research within this rapidly evolving area. Based on these findings, the article finally makes ten predictions for SE in the year 2030.
@@ 1. Introduction
Software engineering (SE) is experiencing a transformation of unprecedented speed and scale, driven by the growing augmentation in software engineering processes and software products with Generative Artificial Intelligence (GenAI) (Abrahão et al., 2025; Pezzè et al., 2024; Ahmed et al., 2025). Such GenAI-augmentation offers unprecedented opportunities for efficiently and effectively developing and operating novel kinds of software systems and applications, but at the same time challenges established principles, practices, tasks, activities and processes of software engineering (Ahmed et al., 2025; Terragni et al., 2024; Banh et al., 2025; Gao et al., 2025). It requires rethinking existing software development lifecycle models (SDLC models), particularly questioning long-held assumptions about the roles of software engineers, the dynamics of hybrid human-AI teams, and the very nature of software artifacts, which now include code, data, and GenAI models.
@@ 2. Methodology
Figure 1. The design science approach for building a roadmap on GenAI-augmented processes and products in SE:
@@ 2.2.1. Awareness - Rapid Literature Reviews
When the research focus is well defined and unambiguous, as in our case, rapid literature reviews (RLRs) constitute pragmatic evidence synthesis methods designed to deliver timely and resource-efficient insights compared to full-fledged Systematic Literature Reviews (SLRs) or Systematic Mapping Studies (SMSs) (Cartaxo et al., 2018). Unlike SLRs and SMSs, RLRs typically relax some methodological requirements, such as triangulation among multiple researchers and systematic validation of study selection and data extraction steps. Instead, these activities are often conducted by one or at most two researchers, which reduces the possibility of cross-checking but allows the review to be completed within a short time frame (usually one or two months). Despite this lower degree of formality, prior work in software engineering has shown that RLRs still provide valuable evidence to support decision-making and efficiently characterize emerging research domains (Cartaxo et al., 2018; Matalonga et al., 2022). We adopted this approach because, after the FSE workshop held at the end of June (see Section 2.1.1), only a few months remained to obtain an initial evidence-based overview of the four fundamental forms of GenAI augmentation in SE. Consequently, small teams of one or two authors conducted focused RLRs, each team addressing one fundamental form, and completed the activity within two months. This pragmatic choice allowed us to align the literature analysis with the design science process, ensuring that each fundamental form was grounded on an initial evidence base despite the short time frame.
@@ 3.1. Structuring of GenAI in SE
As introduced in Section 1, we define the following two dimensions: ”What is augmented by GenAI?”, distinguishing between GenAI-augmented SE processes, meaning that GenAI is used to automate software engineering activities and tasks, and GenAI-augmented software products, meaning that parts of the functionality of a software system or application are not explicitly programmed, but are implemented with GenAI. ”How autonomous is the GenAI augmentation?”, distinguishing between passive roles, meaning that the behavior of GenAI is determined by the methods invoked upon the GenAI model, and active roles, meaning that GenAI has its own thread of control and makes (semi)autonomous decisions about which actions to perform and when (e.g., Agentic AI).
@@
GenAI Copilot: GenAI is used as a tool – either standalone or integrated into an IDE or CI/CD pipeline – to automate various SE tasks. Here, one typically can find solutions from the field of Automated Software Engineering. For the label GenAI Copilot, we took inspiration from Github Copilot (Erhabor et al., 2025) and many other AI/GenAI GenAI Copilots available. Examples of tasks supported by GenAI Copilot include requirement elicitation (Korn et al., 2025), code generation (Jiang et al., 2024), testing (Wang et al., 2024a), and program repair (Zhang et al., 2024).
@@
Note that while we used the 2x2 matrix to define a set of four distinct forms, these may also be considered as four aspects. For example, one may envision that in the future the boundaries between these forms are likely to be fluid and that in a practical setting, multiple forms may be present simultaneously. For instance, a highly advanced GenAI Copilot could start exhibiting proactive behavior, evolving into a GenAI Teammate.
@@ 4.1.1. Review Approach
As this category of GenAI has attracted the most research focus in software engineering to date, we limited our search to top-tier software engineering conferences and journals (ICSE, ASE, FSE, TOSEM, TSE, EMSE, JSS, IST, ASEJ). We restricted the search to the years 2019–present, as the first Transformer-based models (BERT and GPT) were released in 2018, requiring time for adoption by the SE community. To ensure comprehensive coverage, we opted for full-text searches using ACM and IEEE as search engines, as we initially found that many papers only discuss the impact on SE in the body of the paper and not in the title, abstract, or keywords. We narrowed the scope to papers explicitly mentioning the SDLC, development process, or software engineering process in conjunction with LLMs, GenAI, or transformer architecture. These criteria are reflected in the search string presented in Figure 3888The exact search strings can be found in Section A.1.
To further streamline the analysis process, we leveraged Qwen3999https://ollama.com/library/qwen3:8b to analyze the titles, abstracts, and keywords of the retrieved publications. We asked Qwen3 to answer two questions about the publication that reflect the scope of our search: (Q1) Whether the paper explicitly discusses the role and/or implications of GenAI on the SDLC; (Q2) Whether the paper utilizes GenAI or LLMs as a tool to support or automate tasks along the SDLC. Q1 focuses on the publication’s scope, while Q2 excludes GenAIware or GenAI Robot publications, addressed by the following chapters. For each question, Qwen3 provided a yes/no answer along with its reasoning [^10] The exact prompts can be found in Section B.1. The publications positively evaluated by Qwen3 for both questions were then manually inspected by two authors (split equally).
@@ 4.1.2. Review Results
The initial search in ACM and IEEE resulted in 385 publications. Qwen3 classified 161 of these as discussing the role or implications of GenAI on the SDLC and 191 as utilizing GenAI as a tool to support or automate tasks along the SDLC. Of these, 147 met both criteria and were manually inspected. During the manual inspection process, the authors identified that most publications only mention SDLC as motivation or in the related work, but do not describe the implications of GenAI on it (100 articles). Furthermore, many publications made only single statements about the impact or role of GenAI in SDLC, or single tasks of the SDLC that are improved (32 articles). Only a few publications provided a broader discussion of the implications and role of GenAI Copilots on the SDLC (15 papers, see Table 4).
The latter two categories were used to inform our McLuhan’s tetrad (see subsection 4.2), either by deriving insights from the single statements or analyzing the full discussion in the 15 other papers. Furthermore, we included our proposal for the FSE 2025 workshop on “Automated Knowledge Management in the Software Life Cycle” (Keim et al., 2025), as it was one of the starting points for this contribution and fits into the category GenAI Copilot.
@@ 5.1.1. Review Approach
The process for conducting the rapid literature review is depicted in Figure 5.
The used search string is reported in Appendix A.2 and illustrated in Figure 5. The effort bound was set to 200 results; however, in both cases, this limit was never reached. The search was performed using the popular tool Publish or Perish [^11] https://harzing.com/resources/publish-or-perish to facilitate a parameterized search on Google Scholar and Scopus. To include only the relevant sources in the non-peer-reviewed databases, the number of citations ranked the papers, and only those publicly accessible, written in English, with at least one citation were considered for inclusion in the initial pool of sources. Two inclusion criteria were set to ensure the inclusion of papers only relevant to the topic of Generative AI used as an Agent included in the SDLC: (i) the approach presented in the paper had to discuss multiple agents being able to act autonomously or on demand with additional capabilities beyond purely conversational ones (chatbots); (ii) the paper needed to explicitly discuss implication of the approach on the software development lifecycle.
@@ 5.1.2. Review Results
A total of 77 were identified via direct search to be considered for inclusion in the analysis; among these, 27 passed the screening phase by applying duplicate removal, inclusion, and exclusion criteria. Among these, three secondary studies were found, and they were used to complement the information synthesized by directly analyzing primary sources.
@@ 6.1.1. Review Approach
Finally, we manually processed the remaining papers by taking care of removing any remaining papers not relevant to the scope of the review. At this step, we noticed many false positives coming from papers concerning the integration of GenAI Copilot into software development tools. Additionally, we discarded all papers presenting or discussing application-specific solutions, which – despite the fact that they represent a valuable reference for the development of GenAIware – are outside the scope of this review. After this last manual filtering step, we were left with 21 papers.
@@ 6.1.2. Review Results
As mentioned before, memory plays a special role (Lu et al., 2024b). Long-term memory plays a central role in dealing with challenges given by hallucinations, outdated knowledge, and domain-specific gaps, thus the introduction of Retrieval Augmented Generation (RAG)-powered applications for information retrieval (Barnett et al., 2024). These applications enable better reliability and trustworthiness of LLM-powered software. Building these applications implies the choice of the appropriate data store (e.g, vector databases for semantic search, that is, replacing traditional lexical search) and the correct preprocessing of the stored documents (Barnett et al., 2024). User data can be stored and retrieved as well, introducing the need for privacy-related requirements.
@@ 7.1.1. Review Approach
As a second step, we then employed Google’s Gemini 2.5 Flash via its programmatic API to aid us in the analysis process. We prompted Gemini to scrutinize the title, abstract and keywords to answer two specific questions: (Q1) whether the paper actually covers the use of GenAI as a Robot (and not as a GenAI Teammate, for instance), and (Q2) whether the paper explicitly discusses impacts or implications for the SDLC. For each of these questions, we not only asked for a yes/no answer but also required Gemini to provide its rationale for giving the respective answer. The full prompt is provided in Appendix B.4. This resulted in 12 publications for which both answers were yes.
Finally, we manually inspected these 12 publications and eliminated the ones that were out of scope, resulting in 5 relevant publications.
@@ 7.2. McLuhan’s Tetrad
As seen above, there is only little existing work on how GenAI Robots may impact SE, indicating that this is a nascent research area. What this means is that the below McLuhan’s tetrad and roadmap items are necessarily more speculative than those for the other, more established forms. To accommodate the fewer existing work, we also explore more general challenges identified for GenAI Robot in the literature, leading to the following items of the tetrad.
@@ 8.5. Roadmap Part X: Cross-form Aspects
Table 16 describes cross-form challenges and opportunities (X1–X8) that extend beyond individual workflows or deployment contexts. X1 highlights the complexity of coordinating process-level and product-level agents, while X2 stresses the absence of robust benchmarks and evaluation methods for measuring GenAI’s impact in engineering practice. Human factors emerge prominently in X3, where changes to team roles, skills, and dynamics require careful study, and in X4, where unresolved questions of intellectual property, licensing, and liability remain. Broader systemic concerns include sustainability and environmental costs (X5), resilience against adversarial manipulation (X6), and the limited transferability of GenAI methods across domains (X7). Finally, X8 underscores the need for continuous adaptation as software ecosystems evolve. Together, these themes point to a research agenda that blends technical innovation with organizational, legal, and ethical considerations to enable trustworthy GenAI augmentation.

### v2

@@ Abstract.
Abstract. Generative AI (GenAI) is rapidly transforming software engineering (SE) practices, influencing how SE processes are executed, as well as how software systems are developed, operated, and evolved. This paper applies design science research to build a roadmap for GenAI-augmented SE. The process consists of three cycles that incrementally integrate multiple sources of evidence, including collaborative discussions from the FSE 2025 “Software Engineering 2030” workshop, rapid literature reviews, and external feedback sessions involving peers. McLuhan’s tetrads were used as a conceptual instrument to systematically capture the transforming effects of GenAI on SE processes and software products. The resulting roadmap identifies four fundamental forms of GenAI augmentation in SE and systematically characterizes their related research challenges and opportunities. These insights are then consolidated into a set of future research directions. By grounding the roadmap in a rigorous multi-cycle process and cross-validating it among independent author teams and peers, the study provides a transparent and reproducible foundation for analyzing how GenAI affects SE processes, methods and tools, and for framing future research within this rapidly evolving area.
@@ 1. Introduction
Software engineering (SE) is experiencing a transformation of unprecedented speed and scale, driven by the growing augmentation in software engineering processes and software products with Generative Artificial Intelligence (GenAI) (Abrahão et al., 2025; Pezzè et al., 2024; Ahmed et al., 2025). Such GenAI-augmentation offers unprecedented opportunities for efficiently and effectively developing and operating novel kinds of software systems and applications, but at the same time challenges established principles, practices, tasks, activities and processes of software engineering (Ahmed et al., 2025; Terragni et al., 2024; Banh et al., 2025; Gao et al., 2025). It requires rethinking existing software development lifecycle models (SDLC models), particularly questioning long-held assumptions about the roles of software engineers, the dynamics of hybrid human-AI teams, and the very nature of software artifacts, which in addition to code and data, now include GenAI models and prompts.
@@
Overall, we make the following main contributions to SE research:
  • Classification of GenAI Augmentation: We provide a classification of GenAI in SE based on the aforementioned two dimensions, resulting in four distinct forms: GenAI Copilot, GenAIware, GenAI Teammate, and GenAI Robot.
  • Analysis of GenAI Effects: We systematically capture the effects of GenAI on SE for each of the above forms, determining what GenAI enhances, reverses, retrieves from the past, and makes obsolete.
  • Comprehensive Research Roadmap: We distill those findings into a detailed roadmap identifying form-specific and cross-cutting research challenges and opportunities. Key areas include prompt engineering, accountability, hybrid human-AI team dynamics, and the coordination of process-level versus product-level agents.
  • Predictions for 2030: We conclude with ten strategic predictions for how SE may look like in the year 2030. This includes the death of manual coding for routine tasks and the shift from the role of SE developer to AI agent orchestrator.
@@ 2. Methodology
Figure 1. The design science approach for building a roadmap on GenAI-augmented processes and products in SE
The design science process structured into three iterative cycles.
@@ 2.2.1. Awareness - Rapid Literature Reviews
When the research focus is well defined and unambiguous, as in our case, rapid literature reviews (RLRs) constitute pragmatic evidence synthesis methods designed to deliver timely and resource-efficient insights compared to full-fledged Systematic Literature Reviews (SLRs) or Systematic Mapping Studies (SMSs) (Cartaxo et al., 2018). Unlike SLRs and SMSs, RLRs typically relax some methodological requirements, such as triangulation among multiple researchers and systematic validation of study selection and data extraction steps. Instead, these activities are often conducted by one or at most two researchers, which reduces the possibility of cross-checking but allows the review to be completed within a short time frame (usually one or two months). Despite this lower degree of formality, prior work in software engineering has shown that RLRs still provide valuable evidence to support decision-making and efficiently characterize emerging research domains (Cartaxo et al., 2018; Matalonga et al., 2022). We adopted this approach because, after the FSE workshop held at the end of June (see Section 2.1.1), only a few months remained to obtain an initial evidence-based overview of the four fundamental forms of GenAI augmentation in SE. Consequently, small teams of one or two authors conducted focused RLRs, each team addressing one fundamental form, and completed the activity within two months. This pragmatic choice allowed us to align the literature analysis with the design science process, ensuring that each fundamental form was grounded on an initial evidence base despite the short time frame. To provide additional transparency and replicability of the RLR process, we provide open access to the data of our literature reviews (person), 2026).
@@ 2.2.2. Solution - McLuhan’s Tetrads for GenAI augmentation in SE
Image shows the four quadrants of a blank tetrad diagram.
@@ 3.1. Structuring of GenAI in SE
As introduced in Section 1, we define the following two dimensions:
  • ”What is augmented by GenAI?”, distinguishing between GenAI-augmented SE processes (meaning that GenAI is used to automate software engineering activities and tasks) and GenAI-augmented software products (meaning that parts of the functionality of a software system or application are not explicitly programmed, but are implemented with GenAI).
  • ”How autonomous is the GenAI augmentation?”, distinguishing between passive roles (meaning that the human-AI interaction is triggered by the human (Weber, 2024; Hassan et al., 2024) and that the behavior of GenAI is determined by the methods invoked upon the GenAI model) and active roles (meaning that the human-AI interaction is triggered by GenAI and that GenAI has its own thread of control and makes (semi)autonomous decisions about which actions to perform and when; e.g., Agentic AI).
@@
GenAI Copilot: GenAI is used as a tool – either standalone or integrated into an IDE or CI/CD pipeline – to automate various SE tasks. Here, one typically can find solutions from the field of Automated Software Engineering. For the label GenAI Copilot, we took inspiration from Github Copilot (Erhabor et al., 2025) and many other AI/GenAI Copilots available. Examples of tasks supported by GenAI Copilot include requirement elicitation (Korn et al., 2025), code generation (Jiang et al., 2024), testing (Wang et al., 2024a), and program repair (Zhang et al., 2024).
@@
Note that while we used the 2x2 matrix to define a set of four distinct forms, in a concrete software systems more than one form may be present simultaneously. For instance, a highly advanced GenAI Copilot could also exhibit some proactive behavior of a GenAI Teammate.
@@ 4.1.1. Review Approach
Image shows the rapid literature review process to determine relevant publications.
As this category of GenAI has attracted the most research focus in software engineering to date, we limited our search to top-tier software engineering conferences and journals (ICSE, ASE, FSE, TOSEM, TSE, EMSE, JSS, IST, ASEJ). We restricted the search to the years 2019–present, as the first Transformer-based models (BERT and GPT) were released in 2018, requiring time for adoption by the SE community. To ensure comprehensive coverage, we opted for full-text searches using ACM and IEEE as search engines, as we initially found that many papers only discuss the impact on SE in the body of the paper and not in the title, abstract, or keywords. We narrowed the scope to papers explicitly mentioning the SDLC, development process, or software engineering process in conjunction with LLMs, GenAI, or transformer architecture. These criteria are reflected in the search string [^8] The exact search strings can be found in Section A.1 presented in Figure 3.
To further streamline the analysis process, we leveraged Qwen3999https://ollama.com/library/qwen3:8b to analyze the titles, abstracts, and keywords of the retrieved publications. We asked Qwen3 to answer two questions about the publication that reflect the scope of our search: (Q1) Whether the paper explicitly discusses the role and/or implications of GenAI on the SDLC; (Q2) Whether the paper utilizes GenAI or LLMs as a tool to support or automate tasks along the SDLC. Q1 focuses on the publication’s scope, while Q2 excludes GenAIware or GenAI Robot publications, as these are addressed in the chapters below. For each question, Qwen3 provided a yes/no answer along with its reasoning [^10] The exact prompts can be found in Section B.1. The publications positively evaluated by Qwen3 for both questions were then manually inspected by two authors (split equally).
@@ 4.1.2. Review Results
The initial search in ACM and IEEE resulted in 374 publications. Qwen3 classified 161 of these as discussing the role or implications of GenAI on the SDLC and 191 as utilizing GenAI as a tool to support or automate tasks along the SDLC. Of these, 147 met both criteria and were manually inspected. During the manual inspection process, the authors identified that most publications only mention SDLC as motivation or in the related work, but do not describe the implications of GenAI on it (100 articles). Furthermore, many publications made only single statements about the impact or role of GenAI in SDLC, or single tasks of the SDLC that are improved (32 articles). Only a few publications provided a broader discussion of the implications and role of GenAI Copilots on the SDLC (15 papers, see Table 4).
The latter two categories were used to inform our McLuhan’s tetrad (see Section 4.2), either by deriving insights from the single statements or analyzing the full discussion in the 15 other papers. Furthermore, we included our proposal for the FSE 2025 workshop on “Automated Knowledge Management in the Software Life Cycle” (Keim et al., 2025), as it was one of the starting points for this contribution and fits into the category GenAI Copilot.
@@
Image shows the four quadrants of the tetrad diagram resulting from the rapid literature review and analysis.
@@ 5.1.1. Review Approach
Image shows the rapid literature review process to determine relevant publications.
@@
The process for conducting the rapid literature review is depicted in Figure 5. The search string is reported in Appendix A.2 and illustrated in Figure 5. The effort bound was set to 200 results; however, in both cases, this limit was never reached. The search was performed using the popular tool Publish or Perish [^11] https://harzing.com/resources/publish-or-perish to facilitate a parameterized search on Google Scholar and Scopus. To include only the relevant sources in the non-peer-reviewed databases, the number of citations ranked the papers, and only those publicly accessible, written in English, with at least one citation were considered for inclusion in the initial pool of sources. Two inclusion criteria were set to ensure the inclusion of papers only relevant to the topic of Generative AI used as an Agent included in the SDLC: (i) the approach presented in the paper had to discuss multiple agents being able to act autonomously or on demand with additional capabilities beyond purely conversational ones (chatbots); (ii) the paper needed to explicitly discuss implication of the approach on the software development lifecycle.
@@ 5.1.2. Review Results
A total of 77 were identified via direct search to be considered for inclusion in the analysis; among these, 27 passed the screening phase by applying duplicate removal, inclusion, and exclusion criteria (details listed in Table 5). Among these, three secondary studies were found, and they were used to complement the information synthesized by directly analyzing primary sources.
@@ 5.2. McLuhan’s Tetrad
Image shows the four quadrants of the tetrad diagram resulting from the rapid literature review and analysis.
@@ 6.1.1. Review Approach
Image shows the rapid literature review process to determine relevant publications.
@@
We manually processed these LLM-selected papers by taking care of removing any remaining papers not relevant to the scope of the review. At this step, we noticed many false positives coming from papers concerning the integration of GenAI Copilot into software development tools. Additionally, we discarded all papers presenting or discussing application-specific solutions, which – despite the fact that they represent a valuable reference for the development of GenAIware – are outside the scope of this review. After this last manual filtering step, we were left with 21 papers.
@@ 6.1.2. Review Results
As mentioned before, memory plays a special role (Lu et al., 2024b). Long-term memory plays a central role in dealing with challenges given by hallucinations, outdated knowledge, and domain-specific gaps, thus the introduction of Retrieval Augmented Generation (RAG)-powered applications for information retrieval (Barnett et al., 2024). These applications enable better reliability and trustworthiness of LLM-powered software. Building these applications implies the choice of the appropriate data store (e.g., vector databases for semantic search, that is, replacing traditional lexical search) and the correct preprocessing of the stored documents (Barnett et al., 2024). User data can be stored and retrieved as well, introducing the need for privacy-related requirements.
@@ 6.2. McLuhan’s Tetrad
Image shows the four quadrants of the tetrad diagram resulting from the rapid literature review and analysis.
Our rapid literature review informed each quadrant of the tetrad as shown in Figure 8.
@@ 7.1.1. Review Approach
Image shows the rapid literature review process to determine relevant publications.
@@
As a second step, we then employed Google’s Gemini 2.5 Flash via its programmatic API to aid us in the analysis process. We prompted Gemini to scrutinize the title, abstract and keywords to answer two specific questions: (Q1) whether the paper actually covers the use of GenAI as a Robot (and not as a GenAI Teammate, for instance), and (Q2) whether the paper explicitly discusses impacts or implications for the SDLC. For each of these questions, we not only asked for a yes/no answer but also required Gemini to provide its rationale for giving the respective answer. The full prompt is provided in Appendix B.4. This resulted in 11 publications for which both answers were yes.
Finally, we manually inspected these 11 publications and eliminated the ones that were out of scope, resulting in 6 relevant publications.
@@ 7.2. McLuhan’s Tetrad
Image shows the four quadrants of the tetrad diagram resulting from the rapid literature review and analysis.
As seen above, there is only little existing work on how GenAI Robots may impact SE, indicating that this is a nascent research area. What this means is that the below McLuhan’s tetrad and roadmap items are necessarily more speculative than those for the other, more established forms. To accommodate the fewer existing work, we also explore more general challenges identified for GenAI Robot in the literature, leading to the following items of the tetrad, also shown in Figure 10.
@@ 8.5. Roadmap Part X: Cross-form Aspects
Tables 16 and  17 describe cross-form challenges and opportunities (X1–X8) that extend beyond individual workflows or deployment contexts. X1 highlights the complexity of coordinating process-level and product-level agents, while X2 stresses the absence of robust benchmarks and evaluation methods for measuring GenAI’s impact in engineering practice. Human factors emerge prominently in X3, where changes to team roles, skills, and dynamics require careful study, and in X4, where unresolved questions of intellectual property, licensing, and liability remain. Broader systemic concerns include sustainability and environmental costs (X5), resilience against adversarial manipulation (X6), and the limited transferability of GenAI methods across domains (X7). Finally, X8 underscores the need for continuous adaptation as software ecosystems evolve. Together, these themes point to a research agenda that blends technical innovation with organizational, legal, and ethical considerations to enable trustworthy GenAI augmentation.